Cathy is 16 years older than rob. The sum of their ages is 46. How old is each?
Answer by t0hierry(194)

cathy's age is c
rob's age is r
c = r+16
c+r =46
c-r = 16
c+r = 46
2c = 62
c = 31
r = 15